# The Vector Grounding Problem

### v2

@@ 6.2.2 The Argument from Pre-Training
We now turn to a more controversial claim: that pre-training alone can, in certain contexts, establish world-involving functions.
@@
During pre-training, LLMs are explicitly trained to minimise the negative log-likelihood of the next token given previous tokens – an objective limited to the statistical distribution of tokens in the training corpus. [^25] See McCoy et al. (2024) for an approach to studying LLMs, which they call the “the teleological approach”, that takes as its starting point this proximate function. They show that LLM behaviour is strongly shaped by the distributional statistics of the training dataset. As we argue here, however, this is compatible with LLMs also, at least in some cases, fulfilling ultimate, world-involving functions. Sensitivity to distributional patterns can be a means to fulfilling ultimate goals, given all the worldly information embedded in such patterns. At first glance, this appears to be a purely infra-linguistic task with no connection to the world. However, as we have seen in subsection 6.1, the patterns of language use in the training corpus are not arbitrary; they reflect systematic regularities in how humans use language to communicate about the world.
From an information-theoretic perspective, there are vastly more ways to linguistically express world states than there are world states themselves. For next-token prediction, two strategies are possible: direct memorisation of observed sequences, or latent variable factorisation that learns to represent underlying causal factors. The parameter efficiency of the latter strategy relative to the former scales with the ratio of linguistic expressions to world states – typically a very large number. This creates a strong selection pressure for internal states that track latent variables, as they enable more efficient prediction across diverse linguistic contexts.
The Information Bottleneck principle provides a formal framework for understanding this process (Tishby & Zaslavsky 2015). Neural networks learn to extract approximate, minimally sufficient statistics of the input with respect to the output. This involves finding a maximally compressed mapping of the input that preserves as much information as possible about the output. In language models, this connection is even more fundamental. As shown by Delétang et al. (2024), minimising the next-token prediction loss is mathematically equivalent to finding the most efficient way to compress tokenised information. For complex data distributions generated by structured world domains, the most efficient compression involves factorising the problem according to the causal structure that generated the data.
@@
Research in mechanistic interpretability provides concrete evidence that models pre-trained on next-token prediction can develop internal representations with extra-linguistic content at least in some domains. One particularly interesting case is Othello-GPT, a transformer model trained solely to predict moves in Othello game transcripts (Li et al. 2023). Despite receiving no explicit information about Othello’s rules or board structure, this model was found to develop internal representations that encode the current game state in a player-centric manner. Nanda et al. (2023) showed through intervention studies that specific directions in the model’s activation space encode board positions as "Mine," "Yours," or "Empty." When they intervened on these representations by ’flipping’ a tile from ’yours’ to ’mine’ in the model’s activation space, the model’s predictions changed in ways consistent with the altered board state. This suggests not only that the model encodes the board state, but that this encoding plays a causal role in determining its predictions of legal next moves. [^26] For a longer philosophical discussion of causal interventions used to understand interrnal states of Othello-GPT, see Millière & Buckner (forthcoming).
@@
The Othello-GPT example illustrates how selection processes during pre-training can establish ultimate functional roles with extra-linguistic correctness conditions. During training, the model developed specific directions in its activation space that linearly encode the board state from the perspective of the current player. These linear representations emerged because they enabled the model to accurately predict legal moves across diverse game contexts. Their function became to carry information about the current state of the Othello board, with accuracy conditions dependent on the actual game state rather than merely on linguistic patterns. This example is particularly compelling because game transcripts are systematically constrained by the rules of Othello – to predict legal moves effectively, the model developed a linearly decodable internal model that tracks the underlying game state.
@@
The case of LLMs trained on diverse corpora is undoubtedly more complex than Othello-GPT. Games represent a special class of domains with respect to the grounding problem. Unlike open-ended conversation, where success criteria may be fuzzy and context-dependent, games typically have explicit, formalised rules that define success and failure categorically. These rules create a clearly bounded domain with a fixed ontology (game pieces, positions, legal moves), and deterministic success conditions. By contrast, general LLMs encounter messy, open-ended data with few rigid underlying rules. To the extent that language models may acquire internal states with world-involving content, similarly to Othello-GPT with respect to the game world, this presumably only happens in limited domains and contexts, in which next-token prediction on text sequences provides the right kind of learning signal. That said, there is converging evidence from mechanistic interpretability research that LLMs have internal representations that carry information about discourse entities and their relations, and that such representations are causally relevant to the generation of outputs (e.g., Li et al. 2021; see Millière & Buckner forthcoming for a review).
A useful perspective for understanding how pre-training can establish world-involving functional roles is to view it through the lens of meta-learning. Meta-learning – often characterized as ‘learning to learn’ – is the process by which a model improves its learning capabilities through experience across multiple tasks, essentially learning how to learn more efficiently. The pre-training of LLMs can be viewed as a form of meta-learning because it trains the model on a diverse distribution of tasks implicitly contained within the next-token prediction objective. When predicting the next token in various contexts across a large corpus, the model encounters many implicit ‘tasks’, ranging from syntactic completion to factual recall, reasoning, and various domain-specific problems. This defines a bi-level optimisation structure characteristic of meta-learning: an outer loop (the training process itself) that optimises model parameters, and an inner loop (the forward pass during inference) that applies these learned parameters to solve new tasks on the fly (Wu & Varshney 2024). This helps explain why pre-trained LLMs can rapidly adapt to new tasks from just a few examples at inference time, a capacity known as in-context learning (ICL): they have essentially learned optimisation strategies that generalise across the diverse task distribution encountered during pre-training, enabling efficient transfer to novel contexts during inference.
Many different approaches to understanding in-context learning have been proposed. One prominent theory suggests that ICL involves implicitly optimising over latent task objectives. For example, von Oswald et al. (2024) showed that autoregressive transformers trained on sequence prediction tasks develop gradient-based optimisation algorithms in their forward pass. This suggests that when pre-trained LLMs encounter a series of examples in their context window, they can perform an implicit optimisation process over the latent task space. In other words, they can identify the underlying task structure from the examples, and construct an internal objective that captures the task function. When the in-context examples reflect an extra-linguistic task, the implicit objective function defined by the task may also be extra-linguistic. [^27] See, for example, Patel & Pavlick (2022)’s experiment which involves prompting the model in a few-short learning regime to learn mappings from colour terms like ‘cyan’ to points in RGB colour space such as [0, 100, 100]. This is an instance of a world-involving function, having a standard of success and failure directly contingent on how the world is. Optimising this implicit function during the forward pass should dynamically select activation patterns that increase the likelihood of outputs satisfying the world-involving task constraints.
@@
In summary, the argument from pre-training suggests that even without post-training methods like preference tuning, language models may develop internal representations with ultimate world-involving functions through the selective process of next-token prediction over structured text data, at least in some limited domains. This can occur because inducing the latent causal structure of the data provides an informationally efficient solution to the prediction task across diverse contexts. Evidence from mechanistic interpretability studies supports this view at least in toy domains, showing that Generative Pre-Trained Models can develop internal states that carry information about extra-linguistic properties and causally influence outputs. Furthermore, the meta-learning perspective on in-context learning suggests that pre-trained LLMs can implicitly optimise ultimate extra-linguistic task functions at inference time without parameter update.
@@ 7.1 Parameter-identical duplicates
The historical component of our account of referential grounding invites consideration of a classic philosophical objection. The ‘swampman’ thought experiment envisions a perfect duplicate of a human spontaneously formed by random chance (in the original thought experiment, by a lightning strike on a swamp), and thus without any causal history. We can construct an analogous thought experiment for LLMs:
@@
If $M_{1}$ achieves referential grounding on our view, does $M_{2}$ also achieve it? Our answer aligns with the standard response to the swampman scenario: despite being parameter-identical to $M_{1}$, $M_{2}$ would lack referential grounding. First, the internal states of the swamp system, at least at the moment of creation, would lack any causal-informational relations to the world, given that, by hypothesis, the system has no causal history. Moreover, the absence of a selection and learning history means that the internal states of the swamp system lack any functions, including functions to carry information about the world Shea (2018). Consequently, there is no principled way to determine which outcomes count as successes versus failures for a swamp system. $M_{2}$ might produce outputs that appear meaningful to human interpreters, but without the selective history that shaped $M_{1}$, there is no objective basis for treating some of its outputs as ‘correct’ and others as ‘incorrect’. The parameter values themselves, divorced from the process that selected them, cannot ground representational content. [^28] This scenario is somewhat akin to the example, mentioned above, of ants forming the shape of words in the sand by pure chance; or of chimps eventually reproducing a Shakespeare play by randomly hitting keys in a typewriter. Such outputs have no intrinsic meaning, as they were not produced in the appropriate ways, though referentially grounded beings, such as ourselves, may extrinsically impose meaning on them.
It is worth noting, however, that referential grounding could be established relatively quickly if $M_{2}$ were subsequently fine-tuned. Even a brief learning history would begin to establish the relevant world-involving functions Shea (2018). What matters for grounding is not the parameter values per se, but the historical process through which those parameters were selected to perform specific functions.
We can also consider a variant of this thought experiment: [^29] We thank an anonymous referee for raising this scenario.
@@
This scenario differs importantly from the standard swampman case, as $M_{2}$ does have a learning history – it just differs from that of $M_{1}$. The answer to whether $M_{2}$ achieves referential grounding depends on whether one accepts our argument from pre-training in Section 6.2.2.
If one accepts that pre-training alone can, in some cases, establish world-involving functional roles (as we argued), then $M_{2}$ could achieve referential grounding despite skipping the post-training phase. However, the fact that $M_{2}$ happens to have identical parameter values to $M_{1}$ would be a mere coincidence irrelevant to the grounding question. Indeed, its referential grounding would be limited to that achieved via pre-training, while the internal states that drew their grounding from preference tuning in $M_{1}$ would be ungrounded in $M_{2}$.
These thought experiments highlight the crucial role that historical relations play in referential grounding. The internal states of a system, no matter how complex or functionally organised, cannot have representational content without the appropriate history. At the same time, our position avoids an overly restrictive view of what learning histories can establish referential grounding, allowing that different training regimes might enable a system to overcome the Vector Grounding Problem through different causal-historical paths.

### v3

@@ 6.2.2 The Argument from Pre-Training
We now turn to a more controversial claim: that pre-training alone can, under specific conditions, establish world-involving representational functions. The proximate function of an LLM during pre-training is to minimise the negative log-likelihood of the next token given a preceding sequence. [^24] See McCoy2024 for an approach to studying LLMs, which they call the “the teleological approach”, that takes as its starting point this proximate function. They show that LLM behaviour is strongly shaped by the distributional statistics of the training dataset. As we argue here, however, this is compatible with LLMs also, at least in some cases, fulfilling ultimate, world-involving functions. Sensitivity to distributional patterns can be a means to fulfilling ultimate goals, given all the worldly information embedded in such patterns. At first glance, this appears to be a purely intra-linguistic objective, with correctness conditions defined entirely by the statistical distribution of tokens in the training corpus. It is not immediately obvious how a selection process governed by such a criterion could give rise to internal states whose ultimate function is to represent extra-linguistic reality.
To succeed at next-token prediction, two strategies are possible: memorising observed sequences or inducing latent factors responsible for those sequences. For large datasets, the latter strategy is far more efficient (deletangLanguageModelingCompression2024). This can create a selection pressure for internal states that track underlying causal factors, as they enable more efficient prediction across diverse linguistic contexts. Such selection pressure may be enough to endow language models with internal states that stand in mapping relations to states of affairs in the world, sometimes called ‘world models’ (yildirimTaskStructuresWorld2023). However, philosophers of representation have long recognised that standing in mapping relations to states of the world—also known as ‘structural resemblance’—is insufficient for representation. Such mappings are cheap: they can be found between any two sets with the same cardinality. Representational content would be trivial and ubiquitous if structural resemblance were sufficient to secure it (Shea2012-SHEMIR). For internal states that merely map onto extra-linguistic domains to become full-fledged representations, they must have the function of representing those worldly domains and be used by the system to generate outputs precisely because they have this representational function (Shea2018, williams2025structuralcorrespondencesgroundreal).
The concept of meta-learning provides a formal framework for understanding how pre-training could, in principle, establish world-involving representational functions. Meta‑learning, or ‘learning to learn’, is a bi‑level optimisation process with two learning loops: the outer loop trains a model over a range of related tasks, while the inner loop refers to the model’s rapid adaptation when it encounters data from a specific task. Rather than seeking perfect performance on any one training task, the outer loop shapes the model’s parameters so that a small number of gradient steps (the inner loop) can quickly tailor those parameters to a new, previously unseen task from the same distribution. The pre-training of LLMs can be viewed as an implicit form of meta-learning (wuMetaLearningPerspectiveTransformers2024). The outer loop is the standard training process: minimising next-token prediction loss via gradient descent on the model’s parameters over a large corpus. This corpus implicitly contains a massive distribution of distinct tasks—translating, answering questions, summarising, solving problems, playing games, etc. The outer loop thus selects for a single set of parameters that makes the model an effective learner for the myriad tasks embedded in its training data. Wu et al. suggest that the inner loop is the forward pass itself. On their analysis, the fixed parameters learned during the outer loop do not encode a static knowledge base, but rather an on-the-fly optimisation algorithm. For any given input sequence (i.e., a task context), the forward pass through the Transformer’s layers functions as an iterative optimisation process. Each layer takes the token representations from the layer below and performs an update, which Wu et al. show can be mathematically interpreted as an optimisation step that refines the representations to make them progressively more suitable for the final next-token prediction.
This perspective provides a plausible, though speculative, mechanism for the emergence of world-involving representational functions during pre-training. The outer loop’s general, domain-agnostic objective (predict the next token) exerts a selection pressure on the dynamics of the inner-loop optimiser (the forward pass). For specific domains represented in the training data, the most efficient strategy for the inner loop to succeed at its task-specific prediction objective may be to construct and manipulate an internal model of the latent, extra-linguistic state of affairs that generated the data. The proximate, intra-linguistic objective of the outer loop would thereby select for an inner-loop process whose success is conditional on tracking extra-linguistic states. In such cases, the resulting internal states would acquire the ultimate function of representing that extra-linguistic domain. Their function would be to carry information about the world because their having been selected by the learning process is explained by the fact that they did so, thereby enabling more accurate predictions.
The argument’s viability rests on two substantive assumptions. The first is that the meta-learning perspective provides a valid mechanistic account of the Transformer’s forward pass during pre-training. The second is that for certain domains, the general pre-training objective (next-token prediction) creates selection pressures for an inner-loop process whose success is conditional upon tracking extra-linguistic objectives. The argument is most plausible for what we might call formally constrained domains, such as logic puzzles or board games. In such domains, the condition specified in the second assumption is most clearly met. The sequences of tokens in the training data (e.g., game moves) are not generated by the contingent communicative intentions of humans in an open-ended world, but are instead governed by objective, formal rules that constitute the domain itself. The statistical regularities in the data are direct consequences of this underlying formal structure. For a model to succeed at its proximate, intra-linguistic task of predicting valid next tokens, it must become sensitive to these non-accidental regularities. The most robust and efficient strategy for achieving this is to induce an internal model of the domain’s rules and states—the very structure that generates the data. This establishes a non-arbitrary, world-involving standard of correctness, allowing the inner-loop selection process to endow internal states with the ultimate function of representing that extra-linguistic domain.
@@
Research in mechanistic interpretability provides concrete evidence that models pre-trained on next-token prediction can develop internal representations with extra-linguistic content at least in formally constrained domains. One particularly interesting case is Othello-GPT, a transformer model trained solely to predict moves in Othello game transcripts (liEmergentWorldRepresentations2023). Despite receiving no explicit information about Othello’s rules or board structure, this model was found to develop internal representations that encode the current game state in a player-centric manner. nandaEmergentLinearRepresentations2023 showed through intervention studies that specific directions in the model’s activation space encode board positions as ‘Mine’, ‘Yours’, or ‘Empty’. When they intervened on these representations by ‘flipping’ a tile from ‘yours’ to ‘mine’ in the model’s activation space, the model’s predictions changed in ways consistent with the altered board state. This suggests not only that the model encodes the board state, but that this encoding plays a causal role in determining its predictions of legal next moves. [^25] For a longer philosophical discussion of causal interventions used to understand internal states of Othello-GPT, see milliereInterventionistMethodsInterpreting2025.
@@
The Othello-GPT example illustrates how selection processes during pre-training may establish representational functions with extra-linguistic correctness conditions. During training, the model developed specific directions in its activation space that linearly encode the board state from the perspective of the current player. These linear representations plausibly emerged and persisted because they enabled the model to accurately predict legal moves across diverse game contexts. Their function thus became that of carrying information about the current state of the Othello board, with accuracy conditions dependent on the actual game state and the learned rules, rather than merely on linguistic patterns. This example is particularly compelling because game transcripts are systematically constrained by the rules of Othello—to predict legal moves effectively, the model developed a linearly decodable internal model that tracks the underlying game state. [^26] Further research suggests that Othello-GPT may learn many independent, localised decision rules (a “bag of heuristics”) rather than a single systematic algorithm for tracking the board state and determining legal moves (jylin04OthelloGPTLearnedBag2024). For instance, researchers identified specific “board pattern neurons” that activate when a particular local arrangement of pieces makes a specific move legal. This is consistent with our argument; the crucial point is that the internal states guiding such heuristics acquire a world-involving representational function through their selection history. A given “board pattern neuron” was selected by the training process because its activation reliably correlated with a specific, rule-governed state of affairs on the Othello board, and detecting this state of affairs improved the model’s ability to predict a legal next move. The correctness condition for this neuron’s firing is therefore extra-linguistic: it is supplied by the actual state of the Othello board as governed by the game’s objective rules. If the neuron fires when the pattern is not present, or fails to fire when it is, it has malfunctioned with respect to the game of Othello, and this malfunction increases the probability of being penalised by the loss function.
@@
The case of LLMs trained on diverse corpora is undoubtedly more complex than Othello-GPT. Games represent a special class of domains with respect to the grounding problem. Unlike open-ended conversation, where success criteria may be fuzzy and context-dependent, games typically have explicit, formalised rules that define success and failure categorically. These rules create a clearly bounded domain with a fixed ontology (game pieces, positions, legal moves), and deterministic success conditions. By contrast, general LLMs encounter messy, open-ended data with few rigid underlying rules. To the extent that language models may acquire internal states with world-involving content purely through pre-training, similarly to Othello-GPT with respect to the game world, this presumably only happens in formally constrained domains, in which next-token prediction on text sequences provides the right kind of learning signal—provided that the two aforementioned assumptions hold.
@@
In summary, the argument from pre-training suggests that even without post-training methods like preference tuning, language models may develop internal representations with ultimate extra-linguistic representational functions through the selective process of next-token prediction over structured text data, at least in some limited, formally constrained domains. Evidence from mechanistic interpretability studies supports this view at least in simple game domains, showing that generative pre-trained models can develop internal states that carry information about extra-linguistic properties and causally influence outputs.
6.2.3 The Argument from In-Context Learning
After pre-training, LLMs can rapidly adapt to new tasks from just a few examples at inference time, a capacity known as in-context learning (ICL). Through ICL, pre-trained LLMs appear capable of ‘learning’ various tasks on the fly without any update to their internal parameters. These tasks can be intra-linguistic (e.g., translating between languages) or extra-linguistic (e.g., answering questions about world capitals). An intriguing question is whether ICL can be sufficient to achieve referential grounding.
Suppose that we invent a new board game called ‘Chroma’, played by placing coloured tiles on a grid. The game has a clear set of rules: for instance, a blue tile can only be placed adjacent to a red or yellow tile, and scoring depends on creating specific geometric patterns. These rules define the game’s ‘world’—a self-contained domain with its own entities (tiles, grid positions), properties (colours), and causal regularities (legal moves, scoring conditions). Now, imagine we present a pre-trained LLM, which has never encountered Chroma before, with a prompt containing the complete rules of the game, followed by a transcript of the first few moves. We then prompt it to generate the next legal and strategically optimal move. For the LLM to succeed at this task, it cannot rely on merely matching statistical patterns from its pre-training data, as Chroma is a novel game. It must, in some sense, induce and apply the rules of the game without further training. To the extent that an LLM can do this through ICL, we argue that this can be understood as a selection process that establishes world-involving representational functions for the model’s internal representations.
Recent research on ICL suggests that in formal domains, Transformers can learn to use the in-context examples to construct an implicit, task-specific objective function and then perform computations functionally equivalent to gradient descent to find a solution—a process called ‘mesa-optimisation’ (oswaldUncoveringMesaoptimizationAlgorithms2024). [^27] Mesa-optimisation at inference time is analogous to the aforementioned inner-loop optimisation during pre-training. The initial layers of the model parse the context and construct what is effectively a training set within the model’s activation space. Subsequent layers then iteratively refine the internal representations, taking steps to minimise an implicit loss function tailored to the task at hand. During ICL, this mesa-optimisation process within the forward pass constitutes a selection process that provides a transient locus of normativity for the system’s internal states, even without parameter updates. It selects for internal activation patterns because they contribute to satisfying the task function defined in context. In the Chroma example, this task function is plausibly extra-linguistic with respect to the game-world. The objective that would have to be mesa-optimised is not to predict a linguistically plausible next token, but to identify a move that is valid according to the game rules. In other words, the correctness conditions for this task are determined not by the statistics of English, but by the objective state of the Chroma game world.
For example, an activation pattern that correctly represents that “placing a blue tile at C4 is a legal move” will be reinforced and stabilised during the forward pass because it helps to minimise the implicit error against the objective of making a valid move. Conversely, a pattern representing an illegal move will be suppressed. These internal states thereby acquire a transient selected-effects function: their function is to carry information about the state of the Chroma game (e.g., the board configuration, legal moves, strategic opportunities) in service of winning the game. This function, and the normativity it entails, is established on-the-fly, for the duration of solving the specific task presented in context, and will cease to exist once that context is deleted or overwritten.
Whether this actually occurs in LLMs is an open question. The emergence of effective mesa-optimisers is not guaranteed and depends on the properties of the pre-training data distribution (zhengMesaOptimizationAutoregressivelyTrained2024). Nonetheless, the argument from in-context learning provides a speculative but mechanistically plausible route by which an LLM could, in principle, acquire short-lived representations with extra-linguistic content on the fly.
@@ 7.1 Parameter-identical duplicates
The historical component of referential grounding invites consideration of a classic philosophical objection. The ‘swampman’ thought experiment envisions a perfect duplicate of a human spontaneously formed by random chance (in the original thought experiment, by a lightning strike on a swamp), and thus without any causal history. We can construct an analogous thought experiment for LLMs:
@@
If $M_{1}$ achieves referential grounding on our view, does $M_{2}$ also achieve it? Our answer aligns with the standard response to the swampman scenario: despite being parameter-identical to $M_{1}$, $M_{2}$ would lack referential grounding. First, the internal states of the swamp system, at least at the moment of creation, would lack any causal-informational relations to the world, given that, by hypothesis, the system has no causal history. Moreover, the absence of a selection and learning history means that the internal states of the swamp system lack any functions, including functions to carry information about the world (Shea2018). Consequently, there is no principled way to determine which outcomes count as successes or failures for a swamp system. $M_{2}$ might produce outputs that appear meaningful to human interpreters, but without the selection history that shaped $M_{1}$, there is no objective basis for treating some of its outputs as ‘correct’ and others as ‘incorrect’. The parameter values themselves cannot ground representational content.
It is worth noting, however, that referential grounding could be established relatively quickly if $M_{2}$ were subsequently fine-tuned. Even a brief learning history would begin to establish the relevant world-involving functions (Shea2018). What matters for grounding is not the parameter values per se, but the historical process through which those parameters were selected to perform specific functions. [^28] A subtle variant of this case is a Swamp LLM with frozen weights that would nonetheless engage in in-context learning, adapting its outputs based on information provided within a conversational context. One might be tempted to argue that this ‘inner loop’ of adaptation within the forward pass constitutes a sufficient selection process to establish world-involving functions ab initio. We contend that it does not. The capacity for ICL is itself a complex functional property that requires a selection history. In a normally trained model, this capacity is the product of the ‘outer loop’ of pre-training, which selects for internal mechanisms that can efficiently adapt to the diverse tasks and statistical patterns encountered in the training data. The ICL mechanism in such a model thus has a selected-effects function: to enable rapid task adaptation based on contextual cues. A Swamp LLM, by contrast, possesses this mechanism by chance, so the mechanism itself lacks any function. Its apparent adaptation is a form of pseudo-learning—a causal process that mimics genuine, normatively-grounded adaptation but lacks the historical basis for correctness conditions. Without a function to represent correctly, its internal states cannot misrepresent, and therefore cannot be said to represent at all. We are grateful to Harvey Lederman for suggesting this example.
We can also consider a variant of this thought experiment:
@@
This scenario differs importantly from the standard swampman case, as $M_{2}$ does have a learning history—it just differs from that of $M_{1}$. The answer to whether $M_{2}$ achieves referential grounding depends on whether one accepts our argument from pre-training in Section 6.2.2. If one accepts that pre-training alone can, in some cases, establish world-involving representational functions, then $M_{2}$ could achieve referential grounding despite skipping the post-training phase. The fact that $M_{2}$ happens to have identical parameter values to $M_{1}$ would be a mere coincidence irrelevant to the grounding question. Indeed, its referential grounding would be limited to that achieved via pre-training, while the internal states that drew their grounding from preference tuning in $M_{1}$ would be ungrounded in $M_{2}$.
These thought experiments highlight the crucial role that a history of selection plays in referential grounding. The internal states of a system, no matter how complex or functionally organised, cannot have representational content without the appropriate history. At the same time, our position avoids an overly restrictive view of which learning histories can establish referential grounding, allowing that different training regimes might enable a system to overcome the Vector Grounding Problem through different causal-historical paths.